# Do Vision and Language Encoders Represent the World Similarly?

### 10 Jan 2024 v1

@@ Abstract
Aligned text-image encoders such as CLIP have become the de-facto model for vision-language tasks. Furthermore, modality-specific encoders achieve impressive performances in their respective domains. This raises a central question: does an alignment exist between uni-modal vision and language encoders since they fundamentally represent the same physical world? Analyzing the latent spaces structure of vision and language models on image-caption benchmarks using the Centered Kernel Alignment (CKA), we find that the representation spaces of unaligned and aligned encoders are semantically similar. In the absence of statistical similarity in aligned encoders like CLIP, we show that a possible matching of unaligned encoders exists without any training. We frame this as a seeded graph-matching problem exploiting the semantic similarity between graphs and propose two methods - a Fast Quadratic Assignment Problem optimization, and a novel localized CKA metric-based matching/retrieval. We demonstrate the effectiveness of this on several downstream tasks including cross-lingual, cross-domain caption matching and image classification.
footnotetext: Corresponding author: mayugmaniparambil@gmail.com
@@ 1 Introduction
Figure 1: CKA vs data scaling. The CKA score between different vision encoders and text encoders illustrates that vision encoders trained on large enough datasets exhibit similar semantic similarity with an unaligned text encoder, in comparison to the semantic similarity between CLIP’s vision and text encoder (represented by the green star). In this analysis, the text encoder used for all vision encoders is kept constant as All-Roberta-large-v1 [28].
@@
Is this training step necessary? In an ideal scenario, we anticipate an alignment between vision and language encoders as they inherently capture representations of the same physical world. To this end, we employ Centered Kernel Alignment (CKA)  [42, 12, 22], which is known for measuring representation similarity both within and between networks. As shown in Figure 1, we measure the CKA between a variety of unaligned vision and language encoders  [16, 47, 28, 37, 8], on the image-caption pairs of the COCO [27] dataset and show that some have comparable scores to that of aligned encoders like CLIP [40], affirmative of semantic similarities.
@@
  • We present a matching method that seeks to find the permutation of the captions that maximizes the CKA. Hence, We formulate maximizing CKA as a quadratic assignment problem and introduce transformations and normalizations that greatly improve the matching performance.
@@ 4 Proposed Method
Consider a set of $N$ image-caption pairs, $\mathcal{S}=\{({\bm{x}}_{i},{\bm{c}}_{i})\}_{i=1}^{N}$, where ${\bm{x}}_{i}\in{\mathcal{X}}$ and $\mathbf{c}_{i}\in{\mathcal{C}}$ represent the $i$-th image and its corresponding caption, respectively. In this particular example, we are performing caption to image retrieval, but it is applicable for the reverse as well. Let ${\bm{f}}:{\mathcal{X}}\mapsto{\mathbb{R}}^{d_{1}}$ and ${\bm{g}}:{\mathcal{C}}\mapsto{\mathbb{R}}^{d_{2}}$ denote some vision and language encoders respectively. The image-caption pairs are mapped into their corresponding sets of representations ${\mathbf{Z}}=[{\bm{z}}_{1},\ldots,{\bm{z}}_{N}]\in{\mathbb{R}}^{d_{1}\times N}$ and ${\mathbf{H}}=[{\bm{h}}_{1},\ldots,{\bm{h}}_{N}]\in{\mathbb{R}}^{d_{2}\times N}$, where ${\bm{z}}_{i}={\bm{f}}({\bm{x}}_{i})$ and ${\bm{h}}_{i}={\bm{g}}({\bm{c}}_{i})$.
As shown in Table 1, the maximum CKA score is obtained on the ground-truth ordering of the representations $\operatorname{CKA}_{\text{max}}=\operatorname{CKA}({\mathbf{K}}_{{\mathbf{Z}}},{\mathbf{K}}_{{\mathbf{H}}})$, where ${\mathbf{K}}_{{\mathbf{Z}}}$ and ${\mathbf{K}}_{{\mathbf{H}}}$ are the kernels for the image and text representations, defined respectively as ${\mathbf{K}}_{\mathbf{Z}}=k({\mathbf{Z}}^{\top},{\mathbf{Z}})$ and ${\mathbf{K}}_{\mathbf{H}}=k({\mathbf{H}}^{\top},{\mathbf{H}})$. In fact, we find that the CKA is sensitive to the data ordering since it monotonically decreases with random shuffling. This motivates our methodology for finding an optimal permutation of the image data that maximizes the CKA.
@@
To solve the aforementioned optimization problem, we explore two main approaches: the Quadratic Assignment Problem (QAP) algorithm and Local CKA-based retrieval and matching. The QAP algorithm provides a global matching solution, seeking the optimal permutation across the entire test set. On the other hand, Local CKA-based retrieval and matching focuses on aligning images and captions using a localized metric, facilitating retrieval on a more granular level. This approach is more suitable where a single query image is given for a set of captions or vice versa.
@@ 5.3 Downstream Tasks
Table 3: Caption matching and retrieval task performance comparison in cross-domain and in-domain settings. Base samples from COCO are utilized for matching/retrieval tasks on queries from NoCaps (cross-domain) and COCO (in-domain). CLIP-V denotes the vision encoder of CLIP. We use the Large version of all vision encoders. Table A.5 shows the reverse setting.
Method Vision Model NoCaps COCO
Cosine Similarity* CLIP 99.5 99.6 97.1 96.1
Linear regression CLIP-V 29.3 44.7 42.7 59.1
ConvNeXt 19.0 28.5 31.3 46.1
DINOv2 38.1 50.3 45.1 65.4
Relative CLIP-V 61.3 37.6 61.6 41.3
representations ConvNeXt 25.5 17.8 38.6 34.1
DINOv2 46.0 46.4 47.7 52.3
Ours: QAP CLIP-V 67.3 - 72.3 -
ConvNeXt 46.7 - 66.1 -
DINOv2 57.7 - 66.0 -
Ours: Local CKA CLIP-V [40] 65.1 60.5 71.9 69.9
ConvNeXt 43.7 44.4 64.8 65.5
DINOv2 58.7 61.8 64.3 70.5
@@ 5.4 Results
Interestingly, ConvNeXt encoder trained on ImageNet has a classification top1 accuracy improvement of over 14% compared to CLIP and Dinov2 while on the caption matching task DinoV2 and CLIP perform much better.
@@
It is possible to push the performance further by using language-specific sentence encoders and we report these results for a few languages in Sec F of supplementary. This is a practical application of our method as we can now turn a well-trained English CLIP model’s vision encoder into a CLIP model for any low-resource language if a text-only Sentence Transformer trained on that language is available.
@@ 8 Conclusion
In this work, we ask the question, ‘Do vision encoders and language encoders represent the world similarly?’ and study this using CKA and a caption-matching task. We find that well-trained vision encoders on sufficiently large datasets exhibit surprisingly high semantic similarity with language encoders comparable to aligned encoders, irrespective of the training paradigm. Inspired by this, we draw parallels between CKA and the QAP matching objective and use seeded graph matching to align vision and language encoders by maximizing CKA. We also devise a local CKA-based metric to enable retrieval between unaligned vision and language encoders demonstrating a better performance than that of relative representations on cross-domain and cross-lingual caption matching/retrieval tasks, ultimately facilitating zero-shot latent space communication between unaligned encoders.
@@ Appendix A Varying the Number of Samples
In Figure A.1, we show QAP and local CKA matching accuracies and retrieval scores for different number of base samples $M$, keeping the number of query samples $N$ constant at 500. It can be observed that as $M$ increases, accuracy/retrieval scores improve, demonstrating the importance of seed initialization for matching algorithms. Figure A.2 shows the accuracy/retrieval scores as $N$ the number of query samples changes keeping the number of base samples constant at M=320. We see that QAP matching accuracy as local CKA based retrieval scores decrease with increase in $N$, but we still get 70% matching accuracy when $\frac{M}{N}=1$.
@@ Appendix D Mathematical Relationship between Local CKA-based Retrieval and Relative Representations
Appendix E Additional Retrieval Results
@@
Appendix F Additional Cross-Lingual Matching Results

### 22 Mar 2024 v2

@@ Abstract
Aligned text-image encoders such as CLIP have become the de-facto model for vision-language tasks. Furthermore, modality-specific encoders achieve impressive performances in their respective domains. This raises a central question: does an alignment exist between uni-modal vision and language encoders since they fundamentally represent the same physical world? Analyzing the latent spaces structure of vision and language models on image-caption benchmarks using the Centered Kernel Alignment (CKA), we find that the representation spaces of unaligned and aligned encoders are semantically similar. In the absence of statistical similarity in aligned encoders like CLIP, we show that a possible matching of unaligned encoders exists without any training. We frame this as a seeded graph-matching problem exploiting the semantic similarity between graphs and propose two methods - a Fast Quadratic Assignment Problem optimization, and a novel localized CKA metric-based matching/retrieval. We demonstrate the effectiveness of this on several downstream tasks including cross-lingual, cross-domain caption matching and image classification. Code available at github.com/mayug/0-shot-llm-vision.
@@ 1 Introduction
Figure 1: For matching, we calculate the kernels for image and text embeddings and employ QAP-based seeded matching to maximize CKA for obtaining the optimal permutation $\bm{P}$. For retrieval, we append query embeddings to base embeddings and retrieve the best caption that maximizes the local CKA for a query image.
@@
Is this training step necessary? In an ideal scenario, we anticipate an alignment between vision and language encoders as they inherently capture representations of the same physical world. To this end, we employ Centered Kernel Alignment (CKA)  [42, 12, 22], which is known for measuring representation similarity both within and between networks. As shown in Figure 2, we measure the CKA between a variety of unaligned vision and language encoders  [16, 47, 28, 37, 8], on the image-caption pairs of the COCO [27] dataset and observe that some have comparable scores to that of aligned encoders like CLIP [40], affirmative of semantic similarities.
@@
  • We present a matching method that seeks to find the permutation of the captions that maximizes the CKA (see Fig. 1). Hence, We formulate maximizing CKA as a quadratic assignment problem and introduce transformations and normalizations that greatly improve the matching performance.
@@ 4 Proposed Method
Consider a set of $N$ image-caption pairs, $\mathcal{S}=\{({\bm{x}}_{i},{\bm{c}}_{i})\}_{i=1}^{N}$, where ${\bm{x}}_{i}\in{\mathcal{X}}$ and $\mathbf{c}_{i}\in{\mathcal{C}}$ represent the $i$-th image and its corresponding caption, respectively. In this particular example, we are performing caption-to-image retrieval, but it is applicable for the reverse as well. Let ${\bm{f}}:{\mathcal{X}}\mapsto{\mathbb{R}}^{d_{1}}$ and ${\bm{g}}:{\mathcal{C}}\mapsto{\mathbb{R}}^{d_{2}}$ denote some vision and language encoders respectively. The image-caption pairs are mapped into their corresponding sets of representations ${\mathbf{Z}}=[{\bm{z}}_{1},\ldots,{\bm{z}}_{N}]\in{\mathbb{R}}^{d_{1}\times N}$ and ${\mathbf{H}}=[{\bm{h}}_{1},\ldots,{\bm{h}}_{N}]\in{\mathbb{R}}^{d_{2}\times N}$, where ${\bm{z}}_{i}={\bm{f}}({\bm{x}}_{i})$ and ${\bm{h}}_{i}={\bm{g}}({\bm{c}}_{i})$.
As shown in Table 1, the maximum CKA score is obtained on the ground-truth ordering of the representations $\operatorname{CKA}_{\text{max}}=\operatorname{CKA}({\mathbf{K}}_{{\mathbf{Z}}},{\mathbf{K}}_{{\mathbf{H}}})$, where ${\mathbf{K}}_{{\mathbf{Z}}}$ and ${\mathbf{K}}_{{\mathbf{H}}}$ are the kernels for the image and text representations, defined respectively as ${\mathbf{K}}_{\mathbf{Z}}=k({\mathbf{Z}}^{\top},{\mathbf{Z}})$ and ${\mathbf{K}}_{\mathbf{H}}=k({\mathbf{H}}^{\top},{\mathbf{H}})$. We find that the CKA is sensitive to the data ordering. Specifically, we shuffle xof data to obtain wrong matches while keeping the remaining 100-xaligned, measure the CKA on each new data set, and observe that it monotonically decreases with random shuffling. This motivates our methodology for finding an optimal permutation of the image data that maximizes the CKA.
@@
To solve the aforementioned optimization problem, we explore two main approaches (visualized in Fig. 1): the Quadratic Assignment Problem (QAP) algorithm and Local CKA-based retrieval and matching. The QAP algorithm provides a global matching solution, seeking the optimal permutation across the query set considered. On the other hand, Local CKA-based retrieval and matching focuses on aligning images and captions using a localized metric, facilitating retrieval on a more granular level. This approach is more suitable where a single query image is given for a set of captions or vice versa.
@@ 5.4 Results
It is interesting that ConvNeXt encoder trained on ImageNet has a classification top1 accuracy improvement of over 14% compared to CLIP and Dinov2 while on the caption matching task DinoV2 and CLIP perform much better.
@@
It is possible to push the performance further by using language-specific sentence encoders and we report these results for a few languages in Sec I of supplementary. This is a practical application of our method as we can now turn a well-trained English CLIP model’s vision encoder into a CLIP model for any low-resource language if a text-only Sentence Transformer trained on that language is available.
@@ 8 Conclusion
In this work, we ask the question, ‘Do vision encoders and language encoders represent the world similarly?’ and study this using CKA and a caption-matching task. We find that well-trained vision encoders on sufficiently large datasets exhibit surprisingly high semantic similarity with language encoders comparable to aligned encoders, irrespective of the training paradigm. Inspired by this, we draw parallels between CKA and the QAP matching objective and use seeded graph matching to align vision and language encoders by maximizing CKA. We also devise a local CKA-based metric to enable retrieval between unaligned vision and language encoders demonstrating a better performance than that of relative representations on cross-domain and cross-lingual caption matching/retrieval tasks, facilitating zero-shot latent space communication between unaligned encoders.
@@ Appendix A Varying the Number of Samples
In Figure A.1, we show QAP and local CKA matching accuracies and retrieval scores for different number of base samples $M$, keeping the number of query samples $N$ constant at 500. It can be observed that as $M$ increases, accuracy/retrieval scores improve, demonstrating the importance of seed initialization for matching algorithms. Figure A.2 shows the accuracy/retrieval scores as $N$ the number of query samples changes keeping the number of base samples constant at M=320. We see that QAP matching accuracy as local CKA-based retrieval scores decrease with increase in $N$, but we still get 70% matching accuracy when $\frac{M}{N}=1$.
@@ Appendix D Mathematical Relationship between Local CKA-based Retrieval and Relative Representations
Appendix E Other text encoders
Evaluating on COCO with M=320 and N=500, Table A.6 shows that DINOv2-large achieves high QAP accuracy and retrieval performance when combined with different text encoders. This underscores the potential of pairing well-trained sentence and vision encoders for achieving high semantic similarity between image and text embeddings
Table A.6: Comparison of CKA, QAP acc. and local CKA retrieval for different text encoders with DINOv2-large image encoder.
Text Encoder Kernel CKA QAP Acc. Ret @ 5 all-roberta-large-v1 0.690 64.93 77.27 paraphrase-distilroberta-base-v2 0.689 65.07 76.33 paraphrase-mpnet-base-v2 0.695 68.20 81.07 sentence-t5-large 0.660 57.87 69.13 sentence-t5-xxl 0.677 63.40 73.00
Appendix F Simple projection
We trained a 2-layer MLP on frozen DINOv2-large encoder till convergence using CLIP loss and MSE loss. For fair comparison with our setting, we use 320 training and 500 query image-text samples. Results in Table A.8 are averaged over 3 seeds. Notably, QAP matching and local-CKA retrieval excel over projection learning, which demands hyperparameter tuning. In contrast, QAP and local-CKA provide a novel, training-free mechanism to evaluate encoder representational similarity, demonstrating effective latent space communication.
Appendix G Effect of unimodal tasks on alignment
Table A.8 shows using ViT, DETR, DPT, and SegFormer vision encoders for local-CKA and QAP matching on COCO captions (M=320, N=500). ViT is trained on ImageNet-1k (classification), DETR on COCO 2017 (detection), DPT on 1.4M depth images (depth estimation), and SegFormer is fine-tuned on ADE20k (semantic segmentation). Results indicate that classification models exhibit higher semantic similarity to all-roberta-large text encoder in QAP accuracy and local-CKA scores than pixel-level tasks such as object detection, segmentation, and depth estimation.
Table A.7: QAP acc. and Top-5 retrieval scores on COCO.
Table A.8: Unimodal tasks’ effect on image-text alignment.
Method QAP acc Ret @ 5
Proj. + MSE 59.8 73.0
Proj. + CLIP 55.4 68.1
QAP 65.9 -
Local CKA 64.3 76.0
Vision model QAP acc Ret @ 5
ViT 35.3 56.1
DETR 26.5 39.8
DPT 22.7 34.1
Segformer 16.8 33.4
Table A.8: Unimodal tasks’ effect on image-text alignment.
Appendix H Additional Retrieval Results
@@
Appendix I Additional Cross-Lingual Matching Results